### v3

# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

Abstract.
Videos that are shot using commodity hardware such as phones and surveillance cameras record various metadata such as time and location. We encounter such geospatial videos on a daily basis and such videos have been growing in volume significantly. Yet, we do not have data management systems that allow users to interact with such data effectively.
In this paper, we describe Spatialyze, a new framework for end-to-end querying of geospatial videos. Spatialyze comes with a domain-specific language where users can construct geospatial video analytic workflows using a 3-step, declarative, build-filter-observe paradigm. Internally, Spatialyze leverages the declarative nature of such workflows, the temporal-spatial metadata stored with videos, and physical behavior of real-world objects to optimize the execution of workflows. Our results using real-world videos and workflows show that Spatialyze can reduce execution time by up to 5.3$\times$, while maintaining up to 97.1% accuracy compared to unoptimized execution.
{NoHyper}\$*\$\$*\$footnotetext: Both authors contributed equally to this research.
1. Introduction
Geospatial videos are videos in which the location and time that they are shot are known. From surveillance footage, autonomous vehicle (AV) cameras to police body cameras, such videos are prevalent in our daily lives. While the volume of such data has grown tremendously (Wright, 2023; Sparks, 2023), we still lack end-to-end systems that can process and query such data effectively. The rise of machine learning (ML) in recent years has aggravated this issue, with the latest deep learning models capable of carrying out various computer vision tasks, such as object detection (Girshick et al., 2014; Jocher et al., 2022; Girshick, 2015; Redmon and Farhadi, 2018; Ren et al., 2016), multi-objects tracking (Bewley et al., 2016; Du et al., 2023; Wojke et al., 2017; Broström, 2022), image depth estimation (Godard et al., 2019), etc.
All these trends have made geospatial video analytics computationally intensive. For instance, a \$2299 Nvidia T4 GPU takes 34 seconds on average to execute a simple workflow of object detection, tracking, and image depth estimation on a 20-second 12-fps video. Modern autonomous driving datasets, such as nuScenes (Caesar et al., 2019), contain 6000 such videos. Running the workflow mentioned above will take 3 full days to run on the entire dataset.
To make matters worse, the lack of programming frameworks and data management systems for geospatial videos has made it challenging for end users to specify their workflows, let alone running them efficiently. Consider a data journalist writing an article on police misconduct and would like to examine the footage collected on police body cameras to look for specific behavior, e.g., police chasing after a suspect. Given today’s technology, they will either need to watch all the footage themself or string together various ML models for video analytics (Kang et al., 2017; Bastani et al., 2020; Bastani and Madden, 2022; Xu et al., 2022) and video processing libraries (e.g., OpenCV (Bradski, 2000), FFmpeg (Tomar, 2006)) to construct their workflow. Sadly, the former is simply infeasible given the amount of video data collected, while the latter requires deep programming expertise that most users do not possess.
We see an opportunity to build a programming framework that bridges video processing and geospatial video analytics. Specifically, given that users constructing workflows on geospatial videos are typically interested in the time and location where such videos are taken, we should exploit the metadata stored in geospatial videos to optimize such workflows. For instance, a car is at (1810.39, 1221.71, 0.90) in the Boston Seaport map, which is at an intersection, at 2:33:59.112 PM of September 18, 2018. Each object inside a video captures a real-world object; therefore, this object also inherits physical behavior that its real-world counterpart would have had. For example, when users look for a car at a road intersection, we know that a road intersection need to be visible in the video frames that the car of interest is visible; therefore, video frames without a visible road intersection, determined with geospatial metadata, need not be processed with expensive ML functions. We leverage this already existing geospatial metadata, the inherited physical behaviors, and users’ queries to speed up video processing.
To leverage this insight, we present Spatialyze, a system for geospatial video analytics. To make it easy for users to specify their geospatial video workflows, Spatialyze comes with a conceptual data model where users create and ingest videos into a “world,” and users interact with the world by specifying objects (e.g., cars) and scenarios (e.g., cars crossing an intersection) of interest via Spatialyze’s domain-specific language (DSL) embedded in Python called S-Flow. Spatialyze then efficiently executes the workflow by leveraging various spatial-aware optimization techniques that make use of existing geospatial metadata and assumptions based on inherited physical behavior of objects in the videos. For instance, Spatialyze’s Road Visibility Pruner uses road’s visibility as a proxy for objects’ visibility to prune out video frames. The Object Type Pruner then prunes out objects that are not of interest to the users’ workflow. Spatialyze’s 0pt speeds up object 3D location estimation by replacing a computationally expensive ML-based approach with a geometry-based approach. Finally, the Exit Frame Sampler prunes out unnecessary video frames based on the inherited physical behavior of vehicles and traffic rules. Spatialyze’s optimizations are all driven by various geospatial metadata and real-world physical behavior, with a goal to reduce the number of video frames to be processed and ML operations to be invoked, thus speeding up video processing runtime.
As we are unaware of end-to-end geospatial video analytics systems, we evaluate different parts of Spatialyze against state-of-the-art (SOTA) video analytic tools (OTIF (Bastani and Madden, 2022), VIVA (Romero et al., 2022), EVA (Xu et al., 2022)), a geospatial data analytic tool (nuScenes devkit (Caesar et al., 2019)), and an aerial drone video sensing platform (SkyQuery (Bastani et al., 2021)) on different geospatial video workflows. We are up to 7.3$\times$ faster than EVA in geospatial object detection workflow, 1.06-2.28$\times$ faster than OTIF in object tracking, 1.68$\times$ faster than VIVA, 1.18$\times$ faster than SkyQuery in geospatial object tracking workflow, and 117-716$\times$ faster than nuScenes in geospatial data analytics. We also evaluate each of our optimization techniques by executing each of the queries with and without each of the techniques, shown in the results, achieving 2.5-5.3$\times$ speed up with 83.4-97.1% accuracy on 3D object tracking.
In sum, we make following contributions. In § 4, we present a conceptual data model and a DSL (S-Flow) for geospatial video analytics. Using S-Flow, users construct their analytic workflows simply by following our build-filter-observe paradigm. In § 5, we describe the design of Spatialyze, a fast and extensible systems for processing arbitrary-length geospatial videos in S-Flow. In § 6, we develop 4 optimization techniques that leverage geospatial metadata embedded in videos and inherited physical behavior of objects to speed up geospatial video processing in Spatialyze’s geospatial video analytic workflow. In § 7, we implemented our techniques in Spatialyze and evaluated it against other SOTA video processing techniques as well as ablation studies, using real-world videos from nuScenes (Caesar et al., 2019), VIVA (Romero et al., 2022), and SkyQuery (Bastani et al., 2021).
2. Related Work
Geospatial Video Analytics
Recent use cases for querying geospatial video data (autonomous driving (Kim et al., 2021b), surveillance (Griffin and Miller, 2008), traffic analysis (Gloudemans and Work, 2021), and transshipment activities (McDonald et al., 2021; Park et al., 2020)) have driven attention to building geospatial DBMSs for data storage and retrieval. VisualWorldDB (Haynes et al., 2020) allows users to ingest video data from diverse sources and make them queryable as a single multidimensional object. It optimizes query execution by jointly compressing overlapping videos and reusing results from them. Apperception (Ge et al., 2021) provides a Python API for users to input geospatial videos, organizeing them into a four-dimensional data model, and query video data. While Apperception focuses on organizing and retrieving geospatial video data, it falls short in optimizing query execution. In contrast, Spatialyze provides a comprehensive geospatial video analytic workflow interface for declaratively specifying videos of interest. It further introduces new query optimization techniques by leveraging the inherent geospatial properties of videos.
Video Processing
Researchers have proposed techniques to efficiently detect, track, and estimate 3D locations of objects from videos. Monodepth2 (Godard et al., 2019) estimates per-pixel image depth data using only monocular videos. EVA (Xu et al., 2022) identifies, materializes, and reuses expensive user-defined functions (UDF) on videos for exploratory video analytics with multiple refined and reused queries. However, EVA’s optimization techniques are aimed at scenarios where multiple queries use a high degree of overlapping UDFs. OTIF (Bastani and Madden, 2022) is a general-purpose object tracking algorithm. It uses a segmentation proxy model to determine if frames or regions of frames contain objects that need to be processed with an object detector, along with a recurrent reduced-rate tracking method that speed up tracking by reducing frame rate. VIVA (Kang et al., 2022; Romero et al., 2022) is a video analytic system that allows users to specify relationship opportunities of replacing or filtering one ML model with another. It uses these relationships to speed-up video processing time. However, it does not optimize video processing with geospatial metadata by default, which leaves substantial performance gains unexploited, as our experiments show. Complexer-YOLO (Simon et al., 2019) and EagerMOT (Kim et al., 2021a) propose one-stop machine learning models for efficiently tracking 3D objects using LiDAR data, which is not always available compared to cameras and GPS sensors. In addition, 3D sensing data is large. For instance, storing 850 20-second-long scenes of LiDAR sensor data in nuScenes requires 162 GB, while videos of the same scenes only take up 92 GB. Spatialyze overcomes this limitation by not requiring such information to process geospatial workflows, and instead readily available geospatial metadata to optimize user workflows. In addition, one of our optimization techniques estimates objects’ 3D location 192$\times$ faster than the approach that uses Monodepth2 on average. SkyQuery (Bastani et al., 2021) is a drone video sensing platform. It also provides a tool for querying and visualizing sensed aerial drone videos, but does not optimize video query processing.
3. Usage Scenario
In this section, we give an overview of Spatialyze using an example. Suppose a data journalist is interested in investigating AV footage reports, with the goal to find video snippets where other vehicles crash into the drivers’ vehicle at a road intersection.
Naively, the data journalist would watch all the videos individually to identify parts of videos that are relevant. Even if the journalist has programming skills, they still need to manually invoke ML models to process large volumes of geospatial videos, and possibly optimize model inference by reducing the number of frames to be processed. They need to then merge these results with geospatial metadata (e.g., road information) to extract vehicle geospatial data, and write queries to isolate objects of interest and link them to the geospatial objects for a cohesive output. This process demands high coding expertise, is slow, tedious, and error-prone.
Spatialyze streamlines this process. Instead of manually writing code to run ML models and extract geospatial information, journalists can use Spatialyze to easily construct such workflows. With Spatialyze, they can construct an analytic workflow with using S-Flow’s 3-step paradigm as shown in Fig. 1 to: Build a world, by integrating the video data with the geospatial metadata (e.g., road network and camera configurations); Filter the video for parts of interest, by describing relationships between objects (humans, cars, trucks) and their surrounding geospatial environment (lane, intersection); Observe the filtered video parts, in this case, by saving all the filtered video parts into video files for further examination.
Build a world. In line LABEL:line:world, the user creates a world which represents a geospatial virtual environment to be discussed in § 4. Line LABEL:line:road then initializes the world by loading roadNetwork data, in this case, a set of the Boston Seaport’s road segments. Lines LABEL:line:gsvideo-LABEL:line:end-build load videos with their camera configurations (locations, rotations, and intrinsics (Zhang, 2014)) and associate them to create geospatial videos added into the world. Finally, in line LABEL:line:end-build, they have constructed a data-rich world with video cameras, and road segments, all related spatially and temporally.
Filter for the video parts of interest. From the data journalist’s perspective, they can interact with objects and their trackings from each video added to the world through object() with the world initialized. Here, our journalist wants to find video snippets where other vehicles crash into the driver’s vehicle at a road intersection. They do so by adding a filter to the world to only include objects that match the provided predicates: that they are cars or trucks (line LABEL:line:pred-type), are within 50 meters of the camera (line LABEL:line:pred-distance), are at an intersection (line LABEL:line:pred-contain), and are moving towards the camera (in line LABEL:line:pred-heading). After filter is executed at line LABEL:line:pred-heading, the world will contain only cars or trucks on intersections that are moving towards the camera, as desired.
Observe the filtered video parts. After applying the filter, the world has fewer objects compared to when it was initialized. When the journalist calls saveVideos() in line LABEL:line:observe, Spatialyze saves only video snippets with objects in the current world, significantly shortening the total video length. This allows the journalist to focus only on relevant segments with highlighted objects of interest. As shown, S-Flow provides a declarative interface where users do not specify “how” to find the video snippets of interest; instead, they only need to describe “what” their video snippets of interest look like, and Spatialyze will optimize accordingly.
From the users’ perspective, Spatialyze automatically tracks objects from videos, right after each of them is added to the world. Hence, they can filter the world using these tracked objects. Internally, Spatialyze does not execute any part of the workflow until users call saveVideos(). Deferring executions this way allows Spatialyze to analyze the entire workflow for efficient execution.
We next discuss Spatialyze in detail (in § 5) and how it optimizes workflow execution (in § 6) by leveraging video metadata and the physical behavior of real-world objects.
[…]
world.geogConstruct(type=’intersection’) )
[…]
Figure 1. Geospatial Video Analytic Workflow with Spatialyze
4. Constructing Geospatial Video Analytic Workflows
Spatialyze comes with a DSL called S-Flow embedded in python for users to construct their geospatial video data workflow. In this section, we describe S-Flow with Spatialyze’s conceptual data model.
4.1. Conceptual Data Model
Analyzing videos in a general-purpose programming language requires users to manually extract and manipulate object bounding boxes from video frames and join these with geospatial metadata, which is tedious. Hence, we developed Spatialyze’s data model, a higher-level abstraction, to simplify user interaction with geospatial video. This model includes 3 key concepts: 1) World, 2) Geographic Constructs, and 3) Movable Objects, each discussed further below.
4.1.1. World
As discussed in § 3 and in line LABEL:line:world of Fig. 1, users construct a World as a first step of S-Flow. Inspired by VisualWorldDB, a World represents a geospatial virtual environment which encompasses Geographic Constructs and Movable Objects, all coexisting with spatial relationships. These relationships, combined with the S-Flow, enable users to describe the videos of interest effectively.
The World also reflects the physical world, where Movable Objects exhibit realistic behaviors, like cars following lanes and adhering to speed limits. We refer to these phenomena as inherited physical behaviors of Movable Objects. We will discuss how Spatialyze leverage these behaviors to optimize video processing in § 6.
4.1.2. Geographic Constructs
Each World has a set of Geographic Constructs; each Geographic Construct has a spatial property that defines its area, represented as a polygon. It also has non-spatial properties including construct ID and construct type. We model each Geographic Construct as $(cid,\allowbreak type,\allowbreak((x_{0},y_{0}),\allowbreak(x_{1},y_{1}),\allowbreak...))$ where $cid$ is its identifier; $type$ is the construct type (e.g., lane, intersection); and $x_{i},y_{i}$ is a vertex of the polygon that represents the construct.
4.1.3. Movable Objects
Besides static properties like Geographic Constructs, a World may contain multiple Movable Objects. Being movable, they each have spatiotemporal properties, e.g., its location at a given time. They also have non-spatiotemporal properties including object ID and object type that does not change throughout the object’s life-span. For example, a Movable Object with object type of “car” will always be a “car.” Spatialyze models each Movable Object as $(oid,type,((l_{0},r_{0},p_{0},t_{0}),\allowbreak(l_{1},r_{1},p_{1},t_{1}),...))$, where $oid$ is its identifier; $type$ is the object type, such as car, truck, or human. At time $t_{i}$, $l_{i}=(x_{i},y_{i},z_{i})$ is its location; $r_{i}$ is its 3D rotation, defined as a quaternion (Kuipers, 1999); $p_{i}$ is its type-specific properties.
A Movable Object with $type=camera$ represents a camera that captures videos. It has one type-specific property (Zhang, 2014) $p_{i}=(it_{i},)$, where $it_{i}$ is the intrinsic (Anwar, 2022; Zhang, 2014) of the camera at time $t_{i}$. Despite being Movable Objects, this paper distinguishes camera from other types of Movable Objects with the following reason. Movable Objects with $type=camera$ are parts of users’ inputs into the World, while Spatialyze infers Movable Objects with other types from the visual contents captured by the cameras.
We do not include “video” as a part of our conceptual data model. A video itself does not have any geospatial property, so it does not exist in a World. Similarly, none of the objects tracked in the video exists initially in the World. Internally, only after we bind the video with a camera that captures the video, the video gains geospatial properties from the camera and become a part of the camera. Objects tracked in the video also gain their geospatial properties, as a result of the binding. Spatialyze derives the objects’ geospatial locations, by combining their 2 dimensional bounding boxes and camera’s geospatial information. An object that gains geospatial properties then becomes a Movable Object in the World.
4.2. S-Flow
We design S-Flow based on the data model presented in § 4.1, where users interact and manipulate geospatial objects via Geographic Constructs and Movable Objects, rather than video frames. S-Flow’s language constructs consists of functions that operate on World, RoadNetwork, and GeospatialVideo. Users compose such functions to construct geospatial video analytics workflows.
4.2.1. Camera
Lines LABEL:line:gsvideo-LABEL:line:end-build of Fig. 1 defines a Movable Object with $type=camera$ as discussed in in § 4.1.3. Users initialize a Camera from a list, where the $i$-th element of the list is a dictionary with 4 fields: translation, rotation, intrinsic, and timestamp. They correspond to $l_{i}$, $r_{i}$, $it_{i}$, and $t_{i}$, as defined in § 4.1.3.
4.2.2. GeospatialVideo
A GeospatialVideo is a video enhanced with geospatial metadata. A Video (lines LABEL:line:gsvideo-LABEL:line:end-build of Fig. 1) by itself contains a sequence of images where each represents a frame of the video. Such video exists spatially in the World after its Video object is bound to a Camera to obtain its spatial properties, where each frame $i$ is an image taken from the camera at location $l_{i}$, rotation $r_{i}$, intrinsic $it_{i}$, and at time $t_{i}$. The objects in each frame obtains its geospatial properties after its track is inferred from GeospatialVideo. These objects then become Movable Objects as defined in § 4.1.3.
4.2.3. RoadNetwork
A RoadNetwork is a collection of road segments. Each road segment is a Geographic Construct as defined in § 4.1.2. Besides being a Geographic Construct, a road segment stores its segment heading, which indicate traffic headings in that road segment. A road segment have multiple segment headings (curve lane) or no segment heading at all (intersection). Road segments in the RoadNetwork dataset that we use have following road types: roadsection, intersection, lane, and lanegroup. In S-Flow, we initialize a RoadNetwork by loading from a directory containing files of road network from each type. These road network information is accessible and often provided with AV video datasets (Fritsch et al., 2013; Caesar et al., 2019).
4.2.4. World
Each language construct we introduced so far represents users’ data. Bringing the concept of World described in § 4.1.1 into our DSL, users interact with all of their video data through Geographic Construct and Movable Objects that exists in a World. In a Spatialyze workflow, users interact with the World with this 3-step paradigm: build, filter, and observe. First, users build a World by adding videos and geospatial metadata. Once built, users filter the World to only contain their objects of interest. Finally, users observe the remaining objects through annotated videos.
Build. A World is initially empty. Users add RoadNetwork using addGeogContructs(), following by adding GeospatialVideo via addVideo(). At this point of the workflow, the World contains Geographic Constructs, and Cameras. Spatialyze uses YOLOv5 (Jocher et al., 2022) and StrongSORT (Du et al., 2023) for object detection and object tracking by default. Spatialyze also supports using UDFs for object detection or tracking, as shown commented in line LABEL:line:world of Fig. 1.
Filter. Users then construct predicates to filter objects of their interest. As discussed in § 3, Spatialyze has not extracted detections and trackings of objects from the videos yet at this point. Nonetheless, Spatialyze provides an interface for users to interact with these objects as if they are already exist. For instance, object() refers to an arbitrary Movable Object with $type\neq camera$ in the World that is referenced in the predicate. Calling multiple object() gives users multiple arbitrary Movable Objects that appear together in the same video to use in the predicate. For existing information, camera() refers to an arbitrary Camera . Similarly, geoConstruct(type=...) refers to a Geographic Construct of a certain type. Users can construct a predicate involving multiple Movable Objects , a Camera, or Geographic Constructs, using our provided helper functions. For instance, contains(intersection,obj) determines if intersection as a polygon contains obj as a point. headingDiff(obj,cam,between=[0,9]) determines if the difference of the heading of obj and the heading of cam is between 0 and 9 degrees. In addition, we provide more helper functions and an interface for users to define their own helper functions. Finally, users can chain the filter or use boolean operators to combine predicates.
Observe. After filtering their desired Movable Objects, users observe the filtered World, using one of the following 2 observer functions. First, they can observe the World by watching through snippets of the users’ own input videos. Using saveVideos(file, addBoundingBoxes) function, Spatialyze saves all the video snippets that contains the Movable Objects of interest to a file, where addBoundingBoxes is Boolean flag that indicate whether the bounding box of each Movable Objects should be shown. Second, they can observe the World by getting the Movable Object directly. Using getObjects(), Spatialyze returns a list of Movable Objects.
5. Workflow Execution in Spatialyze
Figure 2. Geospatial Video Analytic Workflow Execution.
Figure 3. Processing Plan for the workflow shown in Fig. 1.
From users’ view, Spatialyze simply filters objects in the World after they call filter(). Then, they can observe the filtered objects afterwards, using one of the observer functions described in § 4.2.4. Internally, Spatialyze only performs video processing to detect objects and compute tracks for them as needed given the workflow. For example, in line LABEL:line:pred-heading of Fig. 1, the user filters on objects’ moving directions. Executing this requires computing the trajectory of objects in the given videos to compute their directions. Given an S-Flow workflow, the goal of Spatialyze is to understand the users’ workflow, recognize and execute the video processing operators that are needed for user’s filter predicates, and return the Movable Objects that satisfy users’ filters.
In § 4, we discuss a high-level interface for Spatialyze that allows users to declaratively construct their geospatial video analytic workflows. Users follow the build-filter-observe pattern without knowing when the actual execution happens. Internally, all execution happens when the users observe the world. In this section, we discuss the benefits of deferred execution and how Spatialyze generates execution plans for a workflow.
5.1. Challenges in Workflow Execution
We identify 3 challenges in building a geospatial video analytics system, specifically designing a system that is fast, capable of processing arbitrary-length videos, and extensible. First, processing videos with ML functions is slow; but Spatialyze can leverage its knowledge of the entire workflow in deciding which video frames and ML functions that must be executed. Second, one minute of a 1080p 12fps decoded video is 4.4 GB. Two days of such video can already fill up one GCP VM with the largest available memory size of 11,776 GB as of 2023. Spatialyze must scale with video sizes with limited memory. Finally, Spatialyze must be extensible with future ML functions and optimization techniques. We discuss how we address these challenges in § 5.2 and § 6.
5.2. System Design & Workflow Execution
To minimize execution time, Spatialyze defers ML inferences, which are expensive, until users observe the World. This way, Spatialyze can optimize the entire workflow execution. Once users observe a World, Spatialyze starts executing users’ workflow in 4 stages sequentially: (1) Data Integrator, (2) Video Processor, (3) Movable Objects Query Engine, and (4) Output Composer; as shown in Fig. 2.
5.2.1. Data Integrator
Spatialyze integrates the input data when users Build their World. It processes the Geographic Constructs by creating tables in a geospatial metadata store to store the RoadNetwork and create spatial indexing for every Geographic Construct. Then, it joins each Video and its corresponding Camera frame-by-frame, by their frame number, resulting in a GeospatialVideo.
5.2.2. Video Processor
This stage takes in users’ GeospatialVideo and users’ filter predicates from their workflow’s Filter. It then extracts Movable Objects from the GeospatialVideo. To do so, it first decodes the video to get image frames. It then executes an object detector to extract bounding boxes of objects from each frame. Based on the bounding box, image frame, and Camera, it estimates the 3D location of each object. At the end, it executes the object tracker on the bounding boxes and image frame to get Movable Objects. By default, Spatialyze uses an off-the-shelf algorithm to implement each step. The video decoder is implemented using OpenCV as it is the fastest option available. The object detector is implemented using YOLO (Jocher et al., 2022), a SOTA object detector used by our prior work (Romero et al., 2022; Xu et al., 2022) with high accuracy, low runtime, and capability to detect multiple types of objects. The object 3D location estimator is implemented using Monodepth2 (Godard et al., 2019) as it is the SOTA image depth estimation that only requires an image as an input. The object tracker is implemented using StrongSORT (Du et al., 2023) as it is a fast and online method, suited for processing large videos.
We designed the video processor to stream video frames through all the above algorithms. Each video frame is pipelined through decoding, object detection, 3D location estimation, and object tracking, hence our video processor only needs to keep $O(1)$ video frames and can handle videos of arbitrary size, addressing the second challenge. While stream processing of videos has been used previously (Broström, 2022), applying it to a streaming video processor is not trivial. For example, the original YOLO-StrongSORT object tracker (Broström, 2022) streams video frames. However, if we extended the tracker with our optimization techniques naively, we would have to hard-code them into the tracker’s implementation. This does not scale to future trackers. To solve this problem (and address the third challenge), we implemented our video processor as a plan of streaming operators shown in Fig. 4. Our baseline operators111 The streaming operators internally process videos and geospatial metadata to extract (2D or 3D) object detections and/or tracks. In contrast, users use predicate operators (e.g., contains, distance) to define filter predicates. include: (1) Video Decoder; (2) Object Detector; (3) 3D Location Estimator; and (4) Object Tracker. Specifically, (1) decodes each video to get a sequence of RGB images representing each frame of the video. (2) detects objects in each frame from (1), where each object and its object type and 2D bounding box are returned. AWorld is 3 dimensional where all objects have 3D locations; therefore, (3) estimates the 3D location of each object from (2) based on its Camera. Lastly, (4) tracks objects by associating detections from (3) between frames. We implemented each operator as an iterator function, where each operator takes in streams of per-frame inputs and return a stream of per-frame outputs. As a result, our video processor can process videos of arbitrary size, including those that are larger than the available memory. In addition, we implemented a caching system for the operators. In Fig. 4, frames is an input to both ObjectDetector and ObjectTracker. Instead of computing the stream of frames twice, Spatialyze iterates through frames once and caches the results. For instance, a frame $i$ is decoded when the ObjectDetector processes it. Spatialyze caches the frame $i$ so that the ObjectTracker can process it without having to decode the frame again. Spatialyze keeps track of the number of the operators that will be using a cached result and evicts a frame as soon as all of the operators finished processing it; therefore, Spatialyze only caches a small number of frames at a given time.
Stateless operators like ObjectDetector or Loc3DEstm process each video frame independently. In contrast, ObjectTracker is a stateful operator that keeps track of what objects have been tracked so far at a given frame. To retain states, an ML function can be implemented as a stateful streaming operator using our provided wrapper with minimal changes to the original code of the function.
[…]
Figure 4. Baseline Video Processing Plan
Executing the plan computes 3D tracks of objects in a video. However, streaming each video through all operators is computationally expensive and not always necessary. Based on the users’ filter predicates from when they Filter the World, Spatialyze only includes the necessary operators in the execution plan. For instance, a predicate based on object types like obj.type=="car" requires objects’ types and hence the execution plan will include lines LABEL:line:plan-decode and LABEL:line:plan-detect. Likewise, a predicate with distance or contain functions requires objects’ 3D location and therefore include lines LABEL:line:plan-decode, LABEL:line:plan-detect, and LABEL:line:plan-3d in the plan. Finally, a predicate with a headingDiff function requires objects’ moving directions and include all lines in the plan.
We implemented each optimization technique in § 6 as a streaming operator. Two of the optimization operators execute users’ predicate early on to reduce the number of input (video frames and objects) into later operators. We designed our video processor that constructs plans using streaming operators; therefore, future optimization techniques and video processing functions can be added, addressing the runtime efficiency and extensibility challenges.
5.2.3. Movable Objects Query Engine
This stage takes in users’ filter predicates from their workflow’s Filter and resulting Movable Objects from the video processor. It then filters the Movable Objects with the users’ predicates in 2 steps. First, it streams the Movable Objects into a newly created table in the geospatial metadata store. The table has a spatial index on objects’ tracks and a temporal index on the time period that the objects exist. Then, it translates the user’s filter predicates into an SQL query and execute it in the geospatial metadata store against the Movable Objects, Cameras, and RoadNetwork. When the user’s predicates involve multiple objects (e.g., distance(car1,car2)<5), Spatialyze self-joins the Movable Objects table to find all objects that coexist in a given time period using the temporal index to speed up the join. When the predicates involve spatial relationships between a Movable Object and a RoadNetwork (e.g., contains(road,car)), Spatialyze uses the the spatial index to speed up the joins between the Movable Objects and RoadNetwork tables. We implemented the metadata store using MobilityDB (Zimányi et al., 2020) as it provides necessary spatial and temporal index for movement objects. The index along with its spatiotemporal data type are suited for representing and processing queries on Movable Objects and Geographic Constructs.
5.2.4. Output Composer
This stage composes the Movable Objects query results to a users’ preferred format for them to observe. Calling getObjects, Spatialyze returns a list of Movable Objects. Calling saveVideos, Spatialyze encodes the frames containing the Movable Objects into video files with colored bounding boxes annotation.
Fig. 2 shows the execution stages for the workflow shown in Fig. 1. It contains all stages mentioned above. A user first build the world by adding data. Spatialyze’s Data Integrator creates an indexed table consisting of RoadNetwork and joins the Video and its corresponding Camera. Then, Spatialyze’s Video Processor creates an optimized video processing plan with 4 processing operators and 4 optimization operators (to be discussed in § 6). Parts of the users’ predicates are executed in Road Visibility Pruner and Object Type Pruner. The Movable Objects Query Engine streams the Movable Objects from Video Processor into the geospatial metadata store and filters them with the users’ predicates. Finally, as the workflow asks for videos, Spatialyze formats and saves the frames that have the filtered objects into video snippet files.
6. Video Processing Optimization
SOTA video inference uses ML models to extract the desired information, such as object positions and their trajectories within each video. Running such models is computationally expensive. Spatialyze’s video processor comes with 4 optimization techniques to reduce such runtime bottleneck. To be discussed below, the idea behind our optimization is to leverage the geospatial properties of the videos, and the Movable Objectsinherited physical behaviors. We implement each optimization as a streaming operator in the Video Processing stage in § 5.2.2. Besides the existing 4 processing operators, Spatialyze may place optimization operators in the middle of the video processing plan to reduce the amount of input into later processing operators based on the filter predicates. In § 6.1, Spatialyze places the Road Visibility Pruner after the Video Decoder, when users specify the Geographic Constructs the objects of interest must be in. Doing so reduces the number of input frames for the rest of the operators in the plan. In § 6.2, If users filter on object types, then Spatialyze places the Object Type Pruner right after the Object Detector. It reduces the number of objects that Spatialyze needs to calculate for their 3D locations and tracking. In § 6.3, Spatialyze replaces the ML-based 3D Location Estimation operator with the much faster 0pt, if Spatialyze can assume that the objects of interest are on the ground, such as cars or people. In § 6.4, Spatialyze places the Exit Frame Sampler in between the 3D Location Estimator and the Object Tracker, if the users’ predicates focus on cars. It utilizes the 3D locations of the objects and the user-provided metadata to reduce the number of frames that the Object Tracker processes.
Fig. 3 shows where the video processor executes all the optimization steps. Line LABEL:line:pred-contain in Fig. 1 corresponds to Road Visibility Pruner. Line LABEL:line:pred-type filters by object type, so Spatialyze adds Object Type Pruner. The objects of interest are cars or trucks, so Spatialyze adds Exit Frame Sampler and 0pt. In the following sections, we discuss each optimization in detail.
6.1. Road Visibility Pruner
In our experiments, the runtime of ML models (Jocher et al., 2022; Zhou et al., 2019; Du et al., 2023) in the video processing plan takes 90% of the entire workflow execution runtime, on average. Hence, the goal of the Road Visibility Pruner is to remove frames that do not contain objects of user’s interest to reduce the overall runtime of the video processing plan.
6.1.1. High-Level Concepts
The Road Visibility Pruner uses Geographic Construct’s visibility as a proxy for a Movable Object’s visibility. Specifically, a predicate includes contains(road,obj) and distance(cam,obj)<d means that obj is not visible in the frames where road is not visible within d meters; therefore, Spatialyze prunes out those frames. In Fig. 1, the Road Visibility Pruner partially executes lines LABEL:line:pred-contain where the user searches for “cars within 50 meters at an intersection.” Spatialyze removes all frames that do not contain visible intersection within 50 meters as these frames will not contain any car of users’ interest.
The Road Visibility Pruner only concerns the visibility of Geographic Constructs. Therefore, it uses Geographic Constructs and Camera to decide whether to prune out a video frame. It uses the intrinsic and extrinsic properties of Camera (Zhang, 2014) to determine the viewable area of each of the camera according to the world coordinate system. [^2] World Coordinate System is a 3 dimensional Euclidean space, representing a World. It then determines if the viewable area of each frame includes any Geographic Construct of interest; frames that do not contain such constructs are dropped and will not be processed in the later processing operators. Since the Road Visibility Pruner will be the first processing operator if applied, it can reduce the number of frames that need to be processed by subsequent ML models in the plan, and hence can substantially reduce overall execution time.
6.1.2. Algorithm
The Road Visibility Pruner works on a per-frame basis. For each video frame $i$ shot by a Camera at location $l_{i}$ with rotation $r_{i}$ and camera intrinsic $it_{i}$, the Road Visibility Pruner performs three steps to determine if the frame should be pruned out. First, it computes the 3D viewable space of the Camera at frame $i$. Then, it uses the viewable space to determine Geographic Constructs that are visible to the Camera at frame $i$. Finally, it uses the users’ filter predicates to determine if there is any Geographic Construct of interest visible in the frame. We explain each step below.
First, the Road Visibility Pruner calculates the 3D viewable space of the Camera at frame $i$. The viewable space is a pyramid (Fig. 5 side view). A⃝ is the location of the camera. B⃝, C⃝, D⃝, and E⃝ represent the top-left, top-right, bottom-right, and bottom-left of the frame at $d$ meters in front of the camera. We compute B⃝, C⃝, D⃝, and E⃝ by converting the mentioned 4 corner points of the frame from pixel (2D) to world (3D) coordinates. This is done in 3 steps by converting a point from: 1) Pixel to Camera Coordinate System; 2) Camera to World Coordinate System; and 3) Pixel to World Coordinate System.
Pixel to Camera Coordinate System. The equation for converting a 3D point $(x_{c},y_{c},z_{c})$ from camera coordinate system333 Camera Coordinate System is a 3 dimensional Euclidean space from the perspective of a camera. The origin point is at the camera position. Z-axis and X-axis point toward the front and right of the camera. Y-axis points downward from the camera. to a 2D point $(x_{p},y_{p})$ in pixel coordinate system is:
(1) $\left[\begin{smallmatrix}x_{p}\\ y_{p}\\ 1\end{smallmatrix}\right]=\left[\begin{smallmatrix}f_{x}&s&x_{0}\\ 0&f_{y}&y_{0}\\ 0&0&1\end{smallmatrix}\right]\times\left[\begin{smallmatrix}x_{c}\\ y_{c}\\ z_{c}\end{smallmatrix}\right]\cdot\frac{1}{z_{c}},\text{where }it_{i}=\left[\begin{smallmatrix}f_{x}&s&x_{0}\\ 0&f_{y}&y_{0}\\ 0&0&1\end{smallmatrix}\right]$
where $it_{i}$ represents the camera intrinsic, a set of parameters representing Camera’s focal lengths ($f_{x},f_{y}$), skew coefficient ($s$), and optical center in pixels ($x_{0},y_{0}$). The intrinsic is specific to each camera, independent of its location and rotation. It is used to convert 3D points from camera perspective to 2D points in pixels.
To convert a 2D point in pixel ($x_{p}$ and $y_{p}$), we can find the 3D point in the camera coordinate system ($x_{c}$, $y_{c}$, and $z_{c}$) if we know the distance of the point from the camera ($z_{c}$). From Eq. 1, we have:
(2) $\begin{split}\left[\begin{smallmatrix}x_{c}&y_{c}&z_{c}\end{smallmatrix}\right]^{\top}&=it_{i}^{-1}\times\left[\begin{smallmatrix}x_{p}&y_{p}&1\end{smallmatrix}\right]^{\top}\cdot z_{c}\end{split}$
Modifying Eq. 2 so that the left side is $[x_{c},y_{c},z_{c},1]^{\top}$,
[…]
Camera to World Coordinate System. For this conversion, we set up $R$, a $3\times 3$ matrix, representing the 3D rotation of the camera that is derived from the camera’s rotation quaternion. And, $t$ is a 3D vector that represents the camera’s translation. Then, a $3\times 4$ extrinsic matrix $[R|t]$ converts 3D points from camera coordinate system $(x_{c},y_{c},z_{c})$ to 3D points $(x,y,z)$ in world coordinate system:
(4) $\left[\begin{smallmatrix}x&y&z\end{smallmatrix}\right]^{\top}=[R|t]\times\left[\begin{smallmatrix}x_{c}&y_{c}&z_{c}&1\end{smallmatrix}\right]^{\top}$
Pixel to World Coordinate System. Combining the previous 2 conversions in Eq. 3 and Eq. 4, we can convert 2D points from pixel coordinate system to 3D points in world coordinate system using:
(5) $\begin{split}\left[\begin{smallmatrix}x&y&z\end{smallmatrix}\right]^{\top}&=[R|t]\times C\times\left[\begin{smallmatrix}x_{p}z_{c}&y_{p}z_{c}&z_{c}&1\end{smallmatrix}\right]^{\top}\end{split}$
4 corners of a video frame in World Coordinate System. To convert the 4 corners of a video frame from pixel to world coordinates, we modify Eq. 5 where we substitute $z_{c}$ with $d$, and $x_{p}$ and $y_{p}$ with actual pixel values representing the 4 corners of the video frame; $w$ and $h$ are the width and height in pixels of the video frame, and each of $B,C,D,E$ is a $3\times 1$ vector representing its 3D location.
(6) $\left[\begin{smallmatrix}B&C&D&E\end{smallmatrix}\right]=[R|t]\times C\times\left[\begin{smallmatrix}0&wd&wd&0\\ 0&0&hd&hd\\ d&d&d&d\\ 1&1&1&1\end{smallmatrix}\right]$
After the corners of each frame are translated to world coordinates, the Road Visibility Pruner uses the viewable space to determine the Geographic Constructs that are visible by Camera at frame $i$. Each Geographic Construct’s area is defined using a 2D polygon. Specifically, all polygons representing Geographic Constructs are 2D in the $z=0$ plane. Therefore, the Road Visibility Pruner projects the computed 3D viewable space onto the $z=0$ plane. We define the top-down 2D viewable area as the convex hull (Sklansky, 1982) of the projected vertices A⃝, B⃝, C⃝, D⃝, and E⃝ (yellow highlight in Fig. 5’s top view). We then spatially join the 2D viewable area with Geometric Constructs in the geospatial metadata store, making use of their geospatial indexing created in § 5.2.1. Any Geometric Construct that overlaps with the 2D viewable area is considered visible to the Camera at frame $i$. Once computed, we get visibleGTypes, a set of Geometric Construct types that are visible to the Camera at frame $i$.
Finally, the Road Visibility Pruner uses the users’ filter predicates to determine if there is any Geographic Construct of interest visible in the frame. Examining the contains predicates in the filter, it assigns each contains to $True$ if its first argument is in visibleGTypes and $False$ otherwise. If the value of the transformed filter predicates is $False$, then the Road Visibility Pruner prunes out that frame.
Figure 5. (Side) A 3D viewable space of a camera in pyramid shaped. (Top) Projected area of the pyramid onto $z=0$ plane.
6.1.3. Benefits and Limitations
The Road Visibility Pruner has an insignificant overhead runtime compared to the runtime it can save. Our experiments show that the Road Visibility Pruner takes 0.1% of the video processing runtime to execute, while saving up to 19.6% of the video processing runtime. So in the worst case that Road Visibility Pruner cannot prune out any video frames, the runtime increase is still negligible. However, having the Road Visibility Pruner in the video processing plan may cause tracking accuracy to drop in some scenarios. Using § 3 as an example, if a car starts at an intersection, exits the intersection, and enters an intersection again. During the time that the car is not in any intersection, the camera may not capture any frames containing intersection. The Road Visibility Pruner will prune out the period of video frames where there is no intersection visible to the camera. This can lead to the object tracker considering the car before leaving the intersection and the car after entering the intersection for the second time to be two different cars.
Figure 6. In Fig. 1, the video processor executes expensive ML models only on video frames with a visible intersection.
6.2. Object Type Pruner
SORT-Family algorithms (Du et al., 2023; Wojke et al., 2017; Bewley et al., 2016) use the Hungarian method (Kuhn, 1955) to associate objects between each 2 consecutive video frames. The runtime of the Hungarian method scales with the number of objects to associate in each frame. Reducing the number of objects input into the algorithms reduces the runtime of the object tracker.
6.2.1. High-Level Concepts
Spatialyze only needs to track objects that are of interest as stated in the workflow. The emphObject Type Pruner prunes out irrelevant objects from being tracked. We analyze users’ filter predicates to look for types of objects that are necessary for computing the workflow outputs. Using the predicates in Fig. 1 as an example, the Object Type Pruner partially executes the line LABEL:line:pred-type, where users only look for cars and trucks in the workflow outputs. The trajectories of other objects (e.g., humans) will not appear in the final workflow outputs anyway, so tracking them is unnecessary. Therefore, we only need to input objects that are detected as cars or trucks into object tracker to reduce its workload.
6.2.2. Benefits and Limitations
The Object Type Pruner primarily reduces object tracker’s runtime. Nevertheless, the Object Type Pruner has low overhead runtime of 0.06% and does not require any geospatial metadata, but can reduce up to 69% of the runtime spent on object tracking, which is 26% of the video processor runtime.
6.3. 0pt
Spatialyze uses an object’s distance from camera and its calculated bounding box to estimate its 3D location in our video processing plan. Specifically, we use YOLOv5 (Jocher et al., 2022) to estimate the bounding box and Monodepth2 (Godard et al., 2019) to estimate the object’s distance from camera in our baseline plan. Monodepth2 only requires monocular images (videos) to estimate each object’s distance from camera; therefore, it is extremely flexible in general use cases. Spatialyze can always use Monodepth2 to estimate any object’s distance from camera in a video frame. However, this algorithm is slow and can take upto 48% portion of the total baseline video processing runtime. Therefore, we explored an alternative algorithm that leverage existing geospatial metadata to speed up the process for estimating 3D object location.
6.3.1. High-Level Concepts
0pt leverages Camera to estimate an object’s 3D locations from its 2D bounding box. It is a substitution for Monodepth2 to estimate objects’ 3D locations. 0pt uses basic geometry calculations and requires 2 assumptions to hold. First, the object of interest must be on the ground. We can then assume that the lower horizontal border of the object’s bounding box is the point where the object touches the ground. Second, we can identify the equation of the plane that the object of interest is on. In Spatialyze, the plane $z=0$ is the ground, although our algorithm can be modified to use with any plane. The middle point of the lower border of each object’s 2D bounding box represents the point that it touches the ground (red star in Fig. 7).
6.3.2. Algorithm
0pt performs 2 steps to estimate 3D location of an object from its 2D bounding box. First, based on the second assumption, 0pt finds an equation that represents a Vector of Possible 3D Locations of an object. Second, based on the first and third assumptions, 0pt finds an Exact 3D Location from the vector that intersects with the ground.
Vector of Possible 3D Locations. Using Eq. 5, we convert any 2D point in pixel coordinate system to a 3D point in world coordinate system, if we know $z_{c}$. Since we do not know $z_{c}$ in this case, we can derive a parametric equation representing all possible 3D points (red line in Fig. 7) converted from the 2D point. Let $d$ be the unknown distance of the object from the camera.
(7) $\begin{split}\left[\begin{smallmatrix}x&y&z\end{smallmatrix}\right]^{\top}&=[R|t]\times C\times\left[\begin{smallmatrix}x_{p}d&y_{p}d&d&1\end{smallmatrix}\right]^{\top}\end{split}$
Exact 3D Location. From our assumptions, we know that the bottom line of the object’s bounding box touches the ground. Therefore, $(x,y,z)$ also touches the ground. Because the ground is $z=0$ in our case, we can solve the Eq. 7 for $d$. Once we know $d$, we can solve the parametric Eq. 7 for $x$ and $y$ (blue star in Fig. 7).
6.3.3. Benefits and Limitations
0pt relies on the assumptions mentioned, which do not always hold, especially, the first one. Nevertheless, Spatialyze can analyze users’ filter predicates and tell whether the types of objects of interest can be assumed to touch the ground and only applies the optimization then. For example, “car,” “bicycle,” or “human” touch the ground, while “traffic light” does not. In addition, 0pt is 192$\times$ faster compared than Monodepth2 on average. If one of the estimated 3D locations of an object in a video frame ends up being behind the camera, which means that the object does not touch the ground, Spatialyze can fall back to use Monodepth2 for that frame. Otherwise, 0pt provides 99.48% runtime reduction on average for object 3D location estimation and 52% for the video processing.
Figure 7. 0pt recovers the 3D location (blue star) of a car from its 2D location (red star).
6.4. Exit Frame Sampler
Object trajectories are computed in the object tracker step (§ 5.2.2) using an object tracking algorithm. Such algorithms, including that used in Spatialyze, perform tracking-by-detection (Leal-Taixé, 2014). An object detector identifies objects in video frames but doesn’t link detections across frames to represent the same object. The object tracker associates detections in consecutive frames, creating a chain of associated detections that forms an object trajectory, such as a Movable Object. However, not all detections are necessary in an object trajectory. For instance, a car travels in a straight line from frame 1 to 3, where it is at location (0, 1), (0, 2), (0, 3) at frame 1, 2, 3. The in-between detection (0, 2) at frame 2 is unnecessary as the detections at frame 1 and frame 3 produce the same trajectory.
Spatialyze’s Exit Frame Sampler utilizes users’ geospatial metadata (Camera and RoadNetwork) to heuristically prune out video frames that are likely to contain such unnecessary detections, by sampling only the frames that may contain at least one necessary detection. As a result, the object tracker only needs to perform data association on fewer frames, significantly reducing its runtime.
6.4.1. High-Level Concepts
Suppose a car is visible to a Camera and driving in a lane. As discussed in § 4.1.1, a Movable Object of type “car” has 2 inherited physical behaviors: it follows the lane’s direction and travels at the speed limit regulated by traffic rules. At a particular frame, the Exit Frame Sampler leverages these behaviors to estimate the trajectory of the car so that the object tracker need not track any frame until: i) a car exits its current lane; ii) a car exits the Camera’s view; iii) a new car enters the Camera’s view. These events are sampleEvents. The Exit Frame Sampler samples the frame when a sampleEvent happens for the object tracker to track.
For (i), when the car is in a lane, without object tracking, the Exit Frame Sampler assumes that it follows the lane’s direction until it reaches the end of the lane. After it exits the lane, it may enter an intersection and stop traveling in a straight line. To maintain the accuracy of the trajectory, the object tracker needs to perform data association at the frame right before the car exits the lane. In contrast, if the car is already at an intersection, it may not travel in a straight line; therefore, the object tracker cannot skip any frame.
Event (ii) is the first frame that the car is no longer visible to the Camera. If this event happens before the car exits its lane, although the car’s trajectory still continues in the lane, the object tracker needs to perform data association at the frame before the car is no longer visible to the Camera as the end of the car’s trajectory.
For (iii), if a new car enters the camera’s view, the object tracker needs to start tracking the new car at that frame.
If a frame contains multiple cars, the Exit Frame Sampler skips to the earliest frame that (i), (ii), or (iii) is estimated to happen for any car, which prevents incorrect tracking for any of them. We implement this idea as our sampling algorithm where given a current frame and its detected cars, it samples the next frame for the object tracker to track. This algorithm starting from the first frame of a video to find the next frame and skips to it. Then it repeats the procedure until the end of the video.
Because the algorithm depends on the 2 mentioned inherited physical behaviors, it only works when users filter for Movable Objects with such behaviors, such as vehicles. Therefore, the video processor only adds the Exit Frame Sampler when users’ workflow Filters on the object type being vehicles.
6.4.2. Sampling Algorithm
Fig. 8 shows the pseudocode of each sampleEvent. exitsLane implements (i) when the car exits the lane. Given the car’s location at the current frame (carLoc) and the lane that contains the car, we assume that the car drives in the lane straight along the lane’s direction. In line LABEL:line:exit-lane-loc, we model the car’s movement as a motion tuple (carLoc, lane.direction). We compute the intersection between the car’s motion tuple and the lane as the location where the car exits the lane. Given the car’s speed v, we know the time it reaches the exit location starting from the current frame in line LABEL:line:exit-lane-time. We sample the frame before the car reaches the exit location. exitsCamera implements (ii) when the car exits the camera’s view. Knowing the car’s current location, moving direction, and speed, carMoves computes its location at any given frame in line LABEL:line:car-moves. In line LABEL:line:view-contain, we use the viewable area of the Camera at each frame derived from § 6.1 to find the first frame when the car’s location is not in the camera’s view. This is the frame when the car already exits the camera’s view, so we sample its preceding frame. newCar implements (iii) when a new car enters the Camera’s view. In line LABEL:line:more-car, we sample the next frame that has more cars detected than the current one. The number of cars is a good heuristic to predict that a new car is entering the frame.
def exitsLane(currentFrame, carLoc, lane, v, cameras):
[…]
def exitsCamera(currentFrame, carLoc, lane, v, cameras):
nextFrame = currentFrame + 1
[…]
cameras[nextFrame].time - cameras[currentFrame].time)): nextFrame += 1
return lastFrameBefore(nextFrame)
def newCar(currentFrame):
nextFrame = currentFrame + 1
[…]
return nextFrame
%**** 3-te.tex Line 100 ****
Figure 8. Computing the events that might trigger a frame sample
At the current frame, the Exit Frame Sampler first finds the frame in which we have more detections for sampleEvent (iii). For each car, the Exit Frame Sampler samples frames for for sampleEvents (i) and (ii). Among (i), (ii), (iii), the Exit Frame Sampler choose the earliest sampleEvent as the next frame for the object tracker to track.
6.4.3. Sampling Algorithm Walk-through
Figure 9. Example Results of our Sampling Algorithm
Figure 10. Sample Events exitsLane (i) and exitsCamera (ii)
Frame number 1 labeled as currentFrame in Fig. 9 was taken at 3:51:49.5 PM. Now, we illustrate how the Exit Frame Sampler gets the next frame to sample. In this frame, our object detector only detects one car, $Car_{A}$. The 3D Location Estimator step calculates its 3D location to be (70.92, 74.7, 0). We join the car’s location with the polygons of traffic lanes in our geospatial metadata store to get the geospatial information of traffic lane $Lane_{A}$ that contains the car, specifically, the polygon of $Lane_{A}$ and its direction to be 181° counterclockwise from the east. We speed up this process by using the geospatial index for the polygons. We assume the car’s speed is 25mph based on common traffic rules. We also have the viewable area of all frames.
First, we compute the sampleEvent when there are new cars entering $Camera_{B}$’s view through newCar(currentFrame=1). We find that for the rest of the videos, there is only one car in the frame.
We then compute the sampleEvent when $Car_{A}$ exits $Lane_{A}$ with exitsLane(currentFrame=1,carLoc=(70.92,74.7,0), lane=$Lane_{A}$, v=25mph,$cameras$). In line LABEL:line:exit-lane-loc, $Car_{A}$’s motion tuple intersects the polygon of $Lane_{A}$ at location (66.3, 72.7, 0) and $Car_{A}$ is estimated to reach there at 3:51:49.96 PM in line LABEL:line:exit-lane-time. We then find that the last frame before this time is frame 22, depicted as exitsLane in Fig. 10.
Finally, we compute the sampleEvent when the $Car_{A}$ exits the $Camera_{B}$’s view through exitsCamera(currentFrame=1, carLoc=(70.92,74.7,0),lane=$Lane_{A}$,v=25mph,$cameras$). The camera view does not contain $Car_{A}$’s position at frame 35 computed in line LABEL:line:car-moves. Thus, we sample the frame 34, depicted as exitsCamera in Fig. 10.
Among the 3 sampleEvents, exitsLane in Fig. 10 is the earliest at frame 22. Thus, the Exit Frame Sampler samples the frame before $Car_{A}$ exits its lane at frame 22 labeled as nextFrame in Fig. 9, skipping all frames in between. The Exit Frame Sampler repeats the algorithm to sample frames until it reaches the end of the video.
6.4.4. Benefits and Limitations
To evaluate the efficacy of the Exit Frame Sampler, we define “skip distance” as the number of frames skipped between two non-skipped frames. We quantify tracking performance using the Exit Frame Sampler by summing the runtime of the sampling algorithm and StrongSORT from the current frame to the next unskipped frame, and divided that by the runtime without running the Exit Frame Sampler. We computed the average for all data points with the same skip distance after processing all the videos from the nuScenes dataset. As Fig. 11 shows, the runtime ratio decreases as the skip distance increases. With an average skip of 3.6 frames, per-frame runtime is reduced to 39% of the original.
Meanwhile, skipping too many frames can result in the tracker losing sight of the target vehicle(s). To assess this risk, we computed the F1 score at each skip distance. Here, “prediction” refers to whether an object that appears in frame $f$ is correctly tracked in frame $(f+\textrm{skip distance})$ as predicted by the Exit Frame Sampler. We compared this with the ground truth where each frame is tracked using StrongSORT without the Exit Frame Sampler. Fig. 11 shows that the F1 score remains robust for skip distances under 14, i.e., accuracy does not decrease much within this threshold. With a skip distance of 13, we observed the runtime to be only 28.27% of the original per frame. Hence, we decided to set the maximum skip distance to be 13 to maximize this accuracy vs runtime tradeoff. However, instances with skip distances over 20 have poor accuracy and they are omitted from the graph to focus on effective metrics.
Figure 11. F1 score and runtime reduction for each skip distance when the Exit Frame Sampler can skip at least 1 frame.
7. Evaluation
We have implemented a prototype of Spatialyze and evaluated its components against other state-of-the-art systems. We also evaluate our optimization techniques with an ablation study by comparing Spatialyze’s video processing plans with and without each of them. We explain our experiments setup below.
Queries. We implemented Spatialyze’s workflows using four realistic road scenarios, inspired by queries from Scenic (Kim et al., 2021b), VIVA (Romero et al., 2022) and SkyQuery (Bastani et al., 2021). These workflows test Spatialyze against real-life situations. Each workflow starts by loading RoadNetworks and Videos with their Cameras from each query’s corresponding datasets, similar to lines LABEL:line:world-LABEL:line:end-build in Fig. 1. Then, we filter the videos with one of the queries described below, similar to lines LABEL:line:object-LABEL:line:pred-heading. All queries look for objects closer than 50 meters. Finally, we save the filtered videos into video snippets files, similar to line LABEL:line:observe. Table 1 shows the queries in natural language and S-Flow’s filter predicates.
Q1
A pedestrian at an intersection facing perpendicularly to that of a camera. (vr, s)
contains(intersection, person) & perpendicular(person, camera)
Q2
2 cars at an intersection moving in opposite directions. (vr, s)
contains(intersection, [car1, car2]) & opposite(car1, car2)
Q3
Camera moving in the direction opposite to the lane direction, with another car (which is moving in the direction of the lane) within 10 meters of it. (vr, s)
contains(lane, [camera, car]) & opposite(lane, camera) & sameDirection(lane, car) & (distance(camera, car) < 10)
Q4
A car and a camera moving in the same direction on lanes; 2 other cars moving together on opposite lanes. (vr, s)
contains(lane1, [car1, camera]) & sameDirection(car1, camera) & contains(lane2, [car2, car3]) & sameDirection(car2, car3) & opposite(lane1, lane2)
Q5
A pedestrian is at an intersection. (vr, s)
contains(intersection, person)
Q6
2 cars are at an intersection. (vr, s)
contains(intersection, [car1, car2])
Q7
A car is on a lane within 10 meters of the camera. (vr, s)
contains(lane, camera) & (distance(camera, car) < 10)
Q8
3 cars, each on a lane. (vr, s)
contains(lane, car1) & contains(lane, car2) & contains(lane, car2)
Q9
A car turning left with a pedestrian at an intersection. (vr, vv)
contains(intersection, [car, person]) & turnLeft(car)
Q10
A car stopped in a cycling lane. (vr, sq)
contains(bikeLane, car) & stopped(car)
Table 1. Evaluation query descriptions and their S-Flow predicates (VisualRoad [vr]; Scenic [s]; VIVA [vv]; and SkyQuery [sq]).
Dataset. We use the multi-modal AV datasets from Scenic to evaluate Spatialyze. They comprise of 3 datasets that are needed for executing the 4 queries mentioned: 1) videos from on-vehicle cameras, 2) the cameras’ movement and specification, and 3) road network of Boston Seaport where the videos were shot. The former 2 datasets are directly from nuScenes’s training set (Caesar et al., 2019), and the latter is from Scenic (Kim et al., 2021b; Fremont et al., 2023, 2019). From nuScenes, we randomly sampled 80 out of 467 scenes that are captured at the Boston Seaport from on-vehicle cameras; we chose videos from 3 front cameras for each scene; each video is 20 seconds long at 12 Hz. The runtime of Spatialyze’s execution per video does not depend on the number of input video; therefore, we have determined that 240 sampled videos are sufficient to demonstrate our contributions. To compare Spatialyze with VIVA, in addition to the nuScenes dataset, we use the VIVA-provided Jackson Square Traffic Camera dataset which contains 19,469 5-second-long 1080p videos at 30 Hz of a traffic intersection. Finally, to compare with SkyQuery, we use their aerial drone geospatial video datasets, which contains 17,853 frames of a 1080p top-down video, along with the per-frame GPS information.
Hardware. In our experiments, we use Google Cloud Compute Engine with Nvidia T4; n1-highmem-4 Machine with 2 vCPUs (2 cores), 26 GB of Memory; and Balanced persistent disk.
7.1. Comparison with Competitor Systems
Selecting competitor systems
Since we are not aware of any end-to-end system that focuses on geospatial video analytics, we chose 5 competitor systems to evaluate Spatialyze, where each competitor focuses their novelty of different components covered by Spatialyze.
EVA (Xu et al., 2022). We evaluate Spatialyze against EVA with its SOTA optimization techniques to speed up video processing.
VIVA (Romero et al., 2022). We evaluate Spatialyze against VIVA to emphasize the novelty of our optimization techniques to speed up our end-to-end geospatial video query processing. In addition, the videos, with which VIVA uses to evaluate, are long; we use this fact to evaluate Spatialyze’s capability of processing long videos. Finally, VIVA uses a different ML function (Wojke et al., 2017) to track objects; we evaluate the video processor’s capability of supporting the function.
NuScenes Devkit (Caesar et al., 2019). As the official tool to process NuScenes’ geospatial movement data, we evaluate the runtime-efficiency and the capability of processing large amount of movement objects of the Movable Objects Query Engine against the NuScenes Devkit.
OTIF (Bastani and Madden, 2022). We evaluate Spatialyze against OTIF to emphasize the novelty of our optimization techniques to speed up object tracking.
SkyQuery (Bastani et al., 2021). We evaluate Spatialyze against SkyQuery’s geospatial video query processor. As SkyQuery uses entirely different object detector, tracker, and 3d-estimator, we evaluate the video processor’s ability to support different ML functions and how the existing optimization techniques perform with the new ML functions.
7.1.1. EVA (Xu et al., 2022)
We compare the runtime of the Spatialyze pipeline with that of EVA’s implementation using the same queries. EVA is a VDBMS that allows for the definition of custom user-defined functions (UDFs), and also preforms a number of optimizations. Since EVA executes the queries on a frame-by-frame basis, it cannot compute object tracks (and hence object directions), so we use EVA to evaluate only the object detection capability of Spatialyze with Q5-8. To use of EVA’s caching mechanism, we compare with the runtimes of the queries when they are run in series without resetting EVA and its database (so running Q6 directly after Q5, Q7 directly after Q6, etc). The results are shown in Fig. 12. Our goal in comparing against EVA is to evaluate the efficiency of the two system’s query optimization techniques in § 6.1 and § 6.3.
As shown, Spatialyze performs 2-7.3$\times$ faster than EVA on Q5-7, since Spatialyze does not need to run Monodepth2 on each frame, due to the Object Type Pruner and 0pt. In addition, Spatialyze shaves even more time by employing the Road Visibility Pruner to avoid costly YOLO on frames where they are unnecessary. For Q8, our processing time is comparable to EVA. This is because Spatialyze’s Movable Objects Query Engine finds every possible triplets of cars; thus it performs 2 self-joins internally, while EVA only return video frames with 3 cars or more. We also compare the lines of code (LoC) counts between Spatialyze (14-30) and EVA (58-60), showing that Spatialyze makes query implementation easier and requires less LoC for the same results.
7.1.2. VIVA
We compare the run-time of the Spatialyze workflow with that of VIVA using the traffic application query from VIVA’s evaluation (Q9 in Table 1) as it is VIVA’s only query that involves geospatial content and road network. We run this query against both the VIVA-provided Jackson Square dataset as well as the nuScenes dataset. To match VIVA’s algorithm, we also resize Spatialyze’s input video to a resolution of $360\times 240$, resample the video at 1 FPS, and employ DeepSORT (Wojke et al., 2017) for tracking rather than StrongSORT to demonstrate Spatialyze’s extensibility. We notice that Spatialyze can execute the same workflow at the videos’ native resolution and FPS while VIVA crashes, which demonstrates Spatialyze’s ability to process large videos and the runtime efficiency of the video processor as a result of Spatialyze’s optimization.
Shown in Fig. 12, Spatialyze performs 1.68$\times$ faster compared to VIVA when run on VIVA’s dataset, and 6$\times$ faster when run on the nuScenes dataset. We attribute this increase in performance to Spatialyze’s Object Type Pruner which eliminates the amount of time spent tracking unnecessary objects. VIVA also spends a significantly more time in creating an optimization plan.
In addition, we compare the LoC of the implementations of both VIVA (56) and Spatialyze (51), where Spatialyze provides for a shorter implementation. While we do not have the ground truth to compare with, Spatialyze has higher accuracy in the query results: 45% of tracks returned by Spatialyze are indeed cars turning left. Meanwhile, only 35.7% of the tracks returned by VIVA are correct.
7.1.3. nuScenes
We compare the run-time of the Movable Objects Query Engine stage with the same queries implemented using the nuScenes Devkit (Caesar et al., 2019). NuScenes Devkit is a Python API used to query results from the nuScenes dataset (Caesar et al., 2019). We only compare the Movable Objects Query Engine stage as Devkit operates on already processed videos and ingested object annotations.
The Devkit implementation of Q4 ran out of memory during execution due to the creation of too many Python objects, as a result of materializing all possible combinations of 3 object tuples in memory before filtering them. Shown in Fig. 12, it also takes significantly longer to execute the same query as Spatialyze, with Spatialyze achieving a 117-716$\times$ faster runtime. Two of the main factors that allow Spatialyze to achieve this better execution time are the following. First, Spatialyze pre-processes the road network data, generating additional columns and indexes that allow it to avoid costly joins, which contribute greatly to the large execution time of Devkit. Second, certain Devkit functions perform costly linear algebra that are avoided by Spatialyze using the geospatial metadata store. Comparing the line counts between the different systems, it took 55-123 lines in Devkit to implement the queries while it only took 12-29 lines to do so in S-Flow.
7.1.4. OTIF
We compare the run-time of the video processor of Spatialyze with that of OTIF on object tracking. We run each of the two systems on the sampled nuScenes datasets, and compare the number of frames processed per second for each system. OTIF’s training is 61m37s. After it is trained, OTIF tracks object at 17.3 FPS. For Spatialyze, our video processor applies all optimizations elaborated in § 6, and it processes videos at 18.3-39.5 FPS to get the tracks for each query. Without counting OTIF’s training time, Spatialyze still tracks objects faster.
7.1.5. SkyQuery
SkyQuery is an aerial drone sensing platform for detecting and tracking vehicle from a top-down videos. A user inputs an aerial drone’s video and its geospatial data, then constructs a query to detect and track objects. Spatialyze and SkyQuery both process geospatial videos by 1) detecting objects of interest, 2) finding the objects’ 3D locations, and then 3) tracking the objects.
To compare with SkyQuery, we implemented the 3 steps mentioned as Spatialyze’s video processing steps, and used SkyQuery’s customized YOLOv3 for car detection, their 3D location estimator, and SORT (Bewley et al., 2016) for object tracking. We retained the rest of Spatialyze’s components as well as the optimization steps. Then, we executed both of the systems on Q10 on SkyQuery’s video and geospatial metadata, and compared execution times. As shown in Fig. 12, Spatialyze processes 6.08 FPS, while SkyQuery processes 5.15 FPS, making us 18% faster. We write 15 lines of S-Flow vs 4 lines of SkyQuery to express Q10. The majority of the lines in S-Flow are spent to setup the world and inputting videos and geospatial metadata. While we input videos and geospatial metadata into Spatialyze using S-Flow, we need to specify such input directories in the command line when starting SkyQuery.
As our evaluation uses the same ML models as SkyQuery, our speedup is entirely due to Spatialyze’s leveraging of the query’s semantics to prune video frames and avoid executing ML functions on them. Specifically, Spatialyze uses the Road Visibility Pruner to exclude video frames that do not contain any cycling lanes. We do not evaluate our accuracy in this evaluation as we use the exact same ML inferences functions as those in SkyQuery. As the only optimization technique applied in this evaluation is the Road Visibility Pruner, it only prunes out video frames that do not contain a cycling lane, and thus, does not affect the query result.
Figure 12. Video frames processed per second for each system.
7.2. Ablation Study
To evaluate each of our optimization techniques, we manually construct video processing plans for the following 7 experiment setups. (SB) Baseline is a Baseline plan without any optimization. It includes the 4 processing steps mentioned discussed in § 5.2.2. (S1) is the Baseline plan with the Road Visibility Pruner. (S2) is the Baseline plan with the Object Type Pruner. (S3) is the Baseline plan with Monocular Depth Estimator replaced with the 0pt. (S4) is the Baseline plan with the Exit Frame Sampler. (S5) enables all optimizations except the Exit Frame Sampler. (S6) enables all optimizations. We evaluated Spatialyze on Q1-4 because Spatialyze can apply all or most of the optimization techniques to the executions of such queries; therefore, we can compare the effects of each optimization technique.
7.2.1. Execution time
The average runtime of our workflow execution without any optimization is 34 seconds for each video. When broken down, 0.01% of the time is spent in the Data Integrator; 89.9% in Video Processor; 9.5% in Movable Objects Query Engine; 0.6% in Output Composer. In this section, we evaluate the effect of each of our optimization techniques, by executing Spatialyze on the Q1-Q4. Because the video processing plans for Q3 and Q4 are the same, we do not discuss Q4 in this section.
With all optimization techniques, we achieve 2.5-5.3$\times$ speed up over the baseline plan, as shown in Fig. 13. The Road Visibility Pruner affects the entire plan. It prunes out 3.8% of the frames that do not contain “lane” in Q3 and 21.5% of the frames that do not contain “intersection” in Q1-2, with 3.2-20.0% runtime reduction. The Object Type Pruner only affects StrongSORT. It prunes out 36.5% of objects that are not cars in Q2-4 or trucks and 86.3% that are not pedestrians in Q1, with 7.0-26.3% runtime reduction. The 0pt makes the runtime portion of 3D location estimation insignificant (from 48% to 0.55%). Q1 does not execute the Exit Frame Sampler as it only works with cars or trucks. The Exit Frame Sampler has an overhead runtime of executing § 6.4.2; however, it reduces the number of frames into StrongSORT (13.1%), with 0.8% runtime reduction. We also evaluate the execution time with respect to video length by varying the video length input into the workflow. We found that the execution time of the workflow is linearly proportional to video length and Spatialyze can process videos larger than the available memory.
Figure 13. (Above) Average video processing time for each video of 20 seconds, comparing each optimization technique. (Below) AssA against (SB) vs. other experiment setups.
7.2.2. Accuracy
Our optimization techniques potentially reduce object tracking accuracy as a result of dropping frames. To quantify this, we evaluate our optimization techniques using the execution results from (SB) setup as the ground truth. Then, we compare the execution results from other setups against that of the (SB).
We evaluate our tracking accuracy of each experiment setup by comparing the tracking output from the experiment setup with the tracking output from (SB) setup, using HOTA (Luiten et al., 2020), an evaluation metric for measuring multi-object tracking accuracy. Using the ground truth tracks from a video and prediction tracks from the same video, HOTA returns an Association Accuracy (AssA) score. In this evaluation, the ground truth tracks are the tracking results from the (SB) setup, and the prediction tracks are the tracking results from each of the other setups. We also exclude object detections from the frame pruned by the Road Visibility Pruner because this pruning is a part of users’ predicates and will not be output.
In Fig. 13, the Road Visibility Pruner (S1) causes a small AssA drop of 0.4% for Q3-4 and 4.7% for Q1-2. The Object Type Pruner (S2) also causes a small AssA drop of 2.5% from Q2-4 and 5.3% for Q1. We do not include (S3) into the results since the object tracker only concerns the 2D bounding boxes of objects, 0pt does not affect the AssA of the object tracker. With all 3 optimization techniques (S5), the video processor still produces accurate object tracks of 93.4% on average compared to (SB), while speeding up 2.5-5.3$\times$ of its runtime. With all optimization techniques (S6), the video processor’s accuracy for object tracking is 15.5% on average, while speeding up 1.65-4.32%, additionally.
8. Conclusion
We presented Spatialyze, a system for geospatial video analytics that leverages geospatial metadata and physical properties of objects in videos to speed up video processing. Spatialyze can process videos of arbitrary length by streaming each frame through the operators in workflows expressed using S-Flow. Spatialyze’s DSL allows users to declaratively specify their geospatial video workflows, and our evaluation shows Spatialyze’s efficiency in processing such workflows when compared to SOTA systems.